Cheryl has invested $600 in a mutual fund that charges 6% front-end load. She plans to keep her investment in the fund for the next 10 years. What will be her average annual return, if she expects the fund to earn a 5% reported return in an average year?

4.7%

5%

-1.3%

4.35%
Answer #1

Amount Invested =$600

After front-end load Amount = 600(0.94) = $564

Value after 10 years = 564(1.05)10 = $918.70

So

918.70 = 600(1 + r)10

r = 4.35%